OpenStudy (anonymous):

beginning with the function y=(x-3)^2 +1,what function would represent that same curve shifted down 4 units, to the left 11 units, and opening in a positive direction?

OpenStudy (jdoe0001):

\(\large \begin{array}{cccllll} y=(x&-3)^2 &+1\\ &\uparrow &\uparrow \\ &horizontal\ shift&vertical\ shift\\ &+11&-4\\ &\textit{to the left}&\textit{down} \end{array}\)

OpenStudy (jdoe0001):

recall your transformations section

OpenStudy (anonymous):

a. y=-(x+11)^2 -4 b. y=(x+8)^2 -3 c. y=(x+11)^2 -4 d. y=-(x-7)^2 +20

OpenStudy (jdoe0001):

well, what did you get for the horizontal shifts and vertical shifts, once you apply those values given to the function?

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

y=(x+8)^2-3

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

Wow, thank you you're awesome:)

OpenStudy (jdoe0001):

yw